OpenStudy (anonymous):

How many grams of Na2CO3 are required to make 0.100 moles of Na2CO3? @Abmon98

OpenStudy (abmon98):

Number of Moles=Mass(g)/Molar Mass(g/mol) Look at your periodic table for each element atomic weight in Na2CO3 Na: 23 gram C: 12 gram O: 16 gram Molar Mass(g/mol)=(23*2)+(12*1)+(16*3)=88 gram Number of Moles=Mass(g)/Molar Mass(g/mol) Rearrange your equation Mass(g)=Molar Mass(g/mol)*Number of Moles Mass(g)=0.100*88=8.8

OpenStudy (anonymous):

Thanks For explain And in thatn How many actual molecules of methane are present in 32 grams of methane?

OpenStudy (abmon98):

methane molar mass is 16 12+(1*4)=16 use number of moles=mass(g)/molar mass(g/mol) plug in you given mass 32/16= 2 mole of methane 1 mole consists of 6.02*10^23 molecules (Avogadro constant number) 2 moles consist of x molecules x=2*6.02*10^23/1 molecule

OpenStudy (somy):

mole= mass/Molecular mass u have mass as 32g and molecular mass of Methane u can find from periodic table after you found mole - multiply it by Avogadro number which is \(6.02 \times 10^23\)

OpenStudy (anonymous):

so the answer is 19,264,000 or not?

OpenStudy (somy):

no you calculated it wrong

OpenStudy (somy):

put ( \(6.02 \times 10^23\) ) in brackets like this and the multiply by 2